### v2

Abstract
Set-based state estimation plays a vital role in the safety verification of dynamical systems, which becomes significantly challenging when the system’s sensors are susceptible to cyber-attacks. Existing methods often impose limitations on the attacker’s capabilities, restricting the number of attacked sensors to be strictly less than half of the total number of sensors. This paper proposes a Secure Set-Based State Estimation (S3E) algorithm that addresses this limitation. The S3E algorithm guarantees that the true system state is contained within the estimated set, provided the initialization set encompasses the true initial state and the system is redundantly observable from the set of uncompromised sensors. The algorithm gives the estimated set as a collection of constrained zonotopes, which can be employed as robust certificates for verifying whether the system adheres to safety constraints. Furthermore, we demonstrate that the estimated set remains unaffected by attack signals of sufficiently large and also establish sufficient conditions for attack detection, identification, and filtering. This compels the attacker to inject only stealthy signals of small magnitude to evade detection, thus preserving the accuracy of the estimated set. When a few number of sensors (less than half) can be compromised, we prove that the estimated set remains bounded by a contracting set that converges to a ball whose radius is solely determined by the noise magnitude and is independent of the attack signals. To address the computational complexity of the algorithm, we offer several strategies for complexity-performance trade-offs. The efficacy of the proposed algorithm is illustrated through its application to a three-story building model.
1 Introduction
The interconnectedness of control systems makes them vulnerable to malicious attacks. Sensors are particularly susceptible to these attacks compared to other control system components. The compromised sensor measurements can significantly compromise system safety when employed for control.
To mitigate this risk and ensure accurate state estimation, several algorithms have been proposed recently. Known as resilient or secure state estimation, these algorithms leverage the redundancy of sensors to obtain estimates that converge to a neighborhood of the true state, even in the presence of additive sensor attacks. Secure state estimation algorithms (see, e.g., [14, 11, 19, 29, 16, 3, 9]) are designed to prevent corrupted sensor data from degrading estimation accuracy and maintain reasonable control performance. However, most methods proposed in the literature are point-based, i.e., the estimated state at any given time is a point in the state space. The estimation error bounds of point-based methods are not precise because of their dependence on comparison functions, which are known to be conservative. Moreover, the reliance of point-based methods on majority voting-type algorithms presents a fundamental limitation that only less than half the total number of sensors are allowed to be attacked at any given time. However, an adversary with ample resources would be capable of attacking more than half the number of sensors. While recent efforts investigating the incorporation of intermittent authentication protocols such as cryptography [18] alleviates this fundamental limit, its usage is a tradeoff between computation overhead and estimation accuracy.
1.1 Related Literature
To overcome the limitations of point-based methods and obtain tighter robustness guarantees, the set-based zonotopic filtering paradigm [1, 20, 33, 6, 5, 13, 2, 31] is promising. It has found many real-world applications, including fault diagnosis in industrial systems [7], underwater robotics [17], vehicle localization [8], and leakage detection in water distribution networks [32]. Moreover, in safety-critical applications, guaranteed state inclusion in a bounded set is crucial to verify whether the system adheres to the safety constraints. All these issues further motivate the need for set-based state estimation techniques under unknown system disturbances and measurement errors.
Different from zonotopic estimators, other set-based estimators include interval observers [30, 39], which estimate a box at every time instant, guaranteeing the true state’s inclusion. However, a box fails to capture the inter-dependencies between the state variables and is less accurate [32]. In contrast, zonotopic filters are shown to not only provide an accurate set-based estimate but are also computationally efficient.
Despite the growing importance of secure state estimation in the presence of adversarial attacks on sensors, the existing set-based methods have several shortcomings, including restrictive assumptions, limited robustness to stealthy attacks, and reliance on specific attack strategies. The reachability-based approach of [35] requires the full state vector to be measured by any subset of sensors with cardinality equal to the number of safe sensors. This is a highly restrictive assumption, and without it, the guarantees of estimation accuracy provided in [35] become excessively conservative. Recent works [26, 10, 22] employ standard interval-based or zonotopic filters by identifying and filtering the compromised sensors. However, this approach fails to exclude stealthy attacks on the sensors, which can easily evade the proposed attack detection procedure. As a result, stealthy attacks can significantly affect estimation accuracy by constantly injecting undetectable signals in the sensor measurements, which can accumulate over time and significantly degrade the estimation accuracy. Although this issue is addressed by [24, 23] and [37, 41, 25, 43], the proposed methods are either restricted to a particular attack strategy or rely on the assumption that the attack signal is bounded.
1.2 Our Contribution
In this paper, we propose a novel secure set-based state estimation (S3E) algorithm using zonotopic filtering that does not allow stealthy attacks to significantly corrupt the estimation performance. We also do not assume the boundedness of the attack signals or that the attacker resorts to a particular attack strategy. Moreover, to address the fundamental limitation of point-based secure estimators that strictly less than half the number of sensors can be attacked at any given time, we allow the attacker to compromise all but one sensors under the assumption that the system remains redundantly observable from the remaining uncompromised sensors. Subject to these assumptions, we show that the estimated set obtained from the S3E algorithm is guaranteed to contain the true state of the system.
Although the algorithm presented in this paper can also handle sensor faults, it was developed with adversarial attacks in mind. Attacks differ from faults in that an attack is an intelligently designed signal injected into the sensors to inflict a maximal negative impact on the system and/or evade detection, or in other words, is stealthy. For example, injecting small signals to corrupt the sensor measurements, which cannot be detected at any instant in time, but its effect can accumulate over a long time horizon. Another example is the so-called zero dynamics attack [34], where a specially designed sensor attack signal based on the unstable zero dynamics of the plant can grow unbounded while the plant’s state is driven away from the attack-free trajectory. Such an attack is directed close to the output null space such that the sensor measurement is close to zero (modulo noise). These intelligently designed attacks can cause conventional state estimation algorithms to provide inaccurate state estimates, where the estimation accuracy is dependent on the attack signals. We eliminate the undesirable effect of sensor attacks with the proposed S3E algorithm.
Our algorithm, presented in Section 4, operates in a series of steps at every time instant. Firstly, it calculates the time update by utilizing the model and the bounds on the process noise. Subsequently, subsets of time update consistent with the sensor measurements are computed through a measurement update. The measurement update sets correspond to the possible states in the time update that could have generated the obtained sensor measurements, provided that the exact realization of the measurement noise can be anywhere within the specified bounds described by a zonotope. Some attacked measurement update sets may turn out to be empty if the corresponding attack signal pushes the sensor measurement to a region that is inconsistent with the time update set. In practical scenarios, an attacker could compromise a large subset of sensors. To account for this, we create multiple agreement sets by intersecting various combinations of measurement update sets depending on the parameter of redundant observability of the system. By doing so, we can also eliminate measurement update sets that correspond to attacked sensors. We establish that at least one agreement set contains the true state of the system. Lastly, the estimated set is computed by taking the union of all the agreement sets.
The main strength of our algorithm is that it can handle attacks on different subsets of sensors at any given time. In addition, our approach can accommodate any cardinality of attacked sensors as long as they are less than the total number of sensors. The reason for this is that we do not rely on a simple majority vote among the sensors. Instead, we use a combination of intersections between multiple combinations of measurement update sets, which lead to the agreement sets. These combinations are compared with the model-based time update set in such a way that we can verify the validity of multiple agreement sets even if only one sensor is uncompromised. When at least half of the sensors can be compromised by the attacker, the accuracy of the point-based state estimation algorithms becomes highly susceptible. This is because, in this case, the attacker can easily manipulate the state estimate by injecting arbitrarily large signals. In contrast, our set-based algorithm is resilient to large attack signals as it can naturally detect and identify the corresponding compromised sensors through the measurement update sets or the agreement sets. In Section 5, we provide lower bounds on the magnitude of the attack signals that guarantee attack detection, identification, and/or filtering. Consequently, for the attacker to maintain both stealth and efficacy, they must resort to injecting attack signals of sufficiently small magnitudes.
A shortcoming of our algorithm is that its complexity may increase with time when stealthy attacks are executed, which in our paper are the attack signals with sufficiently small magnitude. Despite this, we argue in Section 5 that it is challenging for the attackers to achieve this worst-case scenario because it requires a complete understanding of both the system and the algorithm, as well as ample computational resources to compute a feasible attack within a single time step. Nonetheless, to tackle this complexity issue, we suggest various strategies to reduce the complexity and facilitate the implementation of the S3E algorithm.
Finally, for analyzing the stability and asymptotic behavior of the S3E algorithm, we integrate a secure point-based state estimator into our algorithm to prune the candidate estimated sets in Section 6. Because of employing a secure point-based estimator, this section relies on the assumption that strictly less than half of the sensors are compromised. However, this assumption is only needed when one requires that the radius of the estimated set is bounded by a contracting ball in a Euclidean space up to a margin that is dependent only on the disturbance and measurement noise. If this assumption is violated, then we can only ensure the boundedness of the radius of the estimated set at every time instant.
In summary, our contributions in this paper include a novel secure set-based state estimation algorithm [^1] The code is available online in our GitHub repository https://github.com/aalanwar/Secure-Set-Based-Estimation. that guarantees the following:
When up to $p-1$ out of total $p$ sensors can be compromised:
  1. 1.
    guaranteed state inclusion, i.e., the true state is always guaranteed to be inside the estimated set;
  2. 2.
    secure state estimation, i.e., large attack signals are automatically discarded, while the impact of stealthy attacks on the estimation accuracy is bounded;
  3. 3.
    sufficient conditions for detection, identification, and filtering of attacks on sensors;
  4. 4.
    methods to address the algorithm’s complexity with minimal compromise on the estimation accuracy.
When strictly less than $p/2$ sensors can be compromised:
  1. (5)
    stability guarantees, i.e., the estimated set remains bounded by a contracting ball whose asymptotic radius only depends on bounds of the process and measurement noise.
To this end, we assume that the system is redundantly observable, i.e., observable from every possible combination of sensors with cardinality less than or equal to the number of safe sensors. This assumption is more realistic than the one in our previous work [28], which required observability from every sensor. Also, the analysis under this assumption turns out to be significantly different and non-trivial compared to [28]. Furthermore, we provide stability guarantees and an attack detection and identification algorithm, which is missing in [28]. Finally, we demonstrate the effectiveness of our proposed method through an illustrative example and a more practical example of a three-story building structure [38] during an earthquake and when an adversary has compromised multiple sensors.
1.3 Outline
The rest of the paper is organized as follows. Section 2 defines the notations and provides the required preliminaries on set representations used in this paper. The main assumptions and the problem are stated in Section 3. Section 4 presents the secure zonotopic state estimation algorithm, provides the inclusion guarantees, and discusses methods to reduce the algorithm’s complexity. Section 5 provides lower bounds on the attacks that can be detected, identified, and/or filtered. Section 6 provides the stability analysis of the estimation algorithm. Finally, Section 7 demonstrates the effectiveness of the proposed algorithm through simulation examples, and Section 8 concludes the paper.
2 Preliminaries
This section establishes the notation used consistently throughout the paper. Additionally, it briefly introduces zonotopes, constrained zonotopes, and their relevant properties and operations employed in the proposed algorithm.
2.1 Notations
The set of real numbers and integers are denoted by $\mathbb{R}$ and $\mathbb{Z}$, respectively. We let $\mathbb{Z}_{\geq i}\coloneq\{i,i+1,i+2,\dots\}$ and $\mathbb{Z}_{[i,j]}\coloneq\{i,i+1,i+2,\dots,j\}$ for $j\geq i$. The Euclidean and maximum norms of a vector $x\in\mathbb{R}^{n}$ are denoted as $\|x\|=\sqrt{x^{\top}x}$ and $\|x\|_{\infty}\coloneq\max_{i\in\{1,\dots,n\}}|x_{i}|$. The vectors of zeros and ones of dimensions $n\times 1$ are denoted as $0_{n}$ and $1_{n}$, respectively. Similarly, $0_{n\times k}$ and $1_{n\times k}$ denote matrices of zeros and ones of dimensions $n\times k$. The identity matrix is $I_{n}\in\mathbb{R}^{n\times n}$. Given a signal $v:\mathbb{Z}_{\geq 0}\rightarrow\mathbb{R}^{n}$, we denote its restriction to the interval $[0,k]$ by $v_{[0,k]}$, for some $k\in\mathbb{Z}_{\geq 0}$. For a finite set $\mathcal{S}$, $|\mathcal{S}|$ denotes its cardinality. The Cartesian product between two compact sets $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ is given by $\mathcal{S}_{1}\times\mathcal{S}_{2}$. The notation (${n\choose k}=\frac{n!}{n!(n-k)!}$ is the number of possible combinations of $k$ out of $n$ elements.
2.2 Set Representations
In this subsection, we introduce the set representations relevant in this paper. Particularly, we define zonotopes and constrained zonotopes, and some set-based operations along with their relevant properties.
2.2.1 Zonotope
Given a center $c_{z}\in\mathbb{R}^{n}$ and a generator $G_{z}\in\mathbb{R}^{n\times\xi_{z}}$, a zonotope $\mathcal{Z}\subset\mathbb{R}^{n}$ is a set
$\mathcal{Z}\coloneq\{c_{z}+G_{z}\beta_{z}:\beta_{z}\in[-1,1]^{\xi_{z}}\}$
where $\xi_{z}$ is the number of generators of $\mathcal{Z}$. That is, a zonotope is an affine transformation of a unit hypercube $\mathcal{H}(0_{n},1)=[-1,1]^{\xi_{z}}$ centered at $0_{n}$ and with radius $1$, where $\xi_{z}\in\mathbb{Z}_{\geq 1}$ is the dimension of the hypercube. Since a zonotope can be completely characterized by its center and generator, the notation $\mathcal{Z}=\langle c_{z},G_{z}\rangle$ is used throughout the paper for brevity.
2.2.2 Basic operations on zonotopes
A matrix $L\in\mathbb{R}^{n^{\prime}\times n}$ multiplied with a zonotope $\mathcal{Z}$ yields a linearly transformed zonotope $L\mathcal{Z}=\langle Lc_{z},LG_{z}\rangle$. Given two zonotopes $\mathcal{Z}_{1}=\langle c_{z_{1}},G_{z_{1}}\rangle$ and $\mathcal{Z}_{2}=\langle c_{z_{2}},G_{z_{2}}\rangle$, each being a subset of $\mathbb{R}^{n}$, their Minkowski sum is given by
$\mathcal{Z}_{1}\oplus\mathcal{Z}_{2}=\langle c_{z_{1}}+c_{z_{2}},[\begin{array}[]{cc}G_{z_{1}}&G_{z_{2}}\end{array}]\rangle.$
Similarly, the Cartesian product of two zonotopes is defined and computed as
$\mathcal{Z}_{1}\times\mathcal{Z}_{2}$ :-$\coloneq\left\{{\left[{\begin{array}[]{c}{{z_{1}}}\\ {{z_{2}}}\end{array}}\right]}:{z_{1}}\in{\mathcal{Z}_{1}},{z_{2}}\in{\mathcal{Z}_{2}}\right\}:$
$=\left\langle\left[\begin{array}[]{c}c_{z_{1}}\\ c_{z_{2}}\end{array}\right],\left[\begin{array}[]{cc}G_{z_{1}}&0\\ 0&G_{z_{2}}\end{array}\right]\right\rangle.$
2.2.3 Constrained zonotope
A constrained zonotope is a set
$\mathcal{Z}\coloneq\{c_{z}+G_{z}\beta_{z}:\beta_{z}\in[-1,1]^{\xi_{z}},A_{z}\beta_{z}=b_{z}\}$
where $c_{z}\in\mathbb{R}^{n}$, $G_{z}\in\mathbb{R}^{n\times\xi_{z}}$, $A_{z}\in\mathbb{R}^{m\times\xi_{z}}$ and $b_{z}\in\mathbb{R}^{m}$ with $m,n\in\mathbb{Z}_{\geq 1}$. That is, a constrained zonotope is an affine transformation of the linearly constrained unit hypercube $\{\beta_{z}\in[-1,1]^{\xi_{z}}:A_{z}\beta_{z}=b_{z}\}$. We use the notation $\mathcal{Z}=\langle c_{z},G_{z},A_{z},b_{z}\rangle$ for a constrained zonotope.
2.2.4 Generalized intersection
Given two constrained zonotopes $\mathcal{Z}\subset\mathbb{R}^{n_{z}}$ and $\mathcal{Y}\subset\mathbb{R}^{n_{y}}$, and a matrix $M\in\mathbb{R}^{n_{y}\times n_{z}}$, the generalized intersection is defined as
$\mathcal{Z}\cap_{M}\mathcal{Y}:=\{z\in\mathcal{Z}:Mz\in\mathcal{Y}\}.$ (1)
Let $\mathcal{Z}=\langle c_{z},G_{z},A_{z},b_{z}\rangle$ with $c_{z}\in\mathbb{R}^{n_{z}}$, $G_{z}\in\mathbb{R}^{n_{z}\times\xi_{z}}$, $A_{z}\in\mathbb{R}^{m_{z}\times\xi_{z}}$, and $b_{z}\in\mathbb{R}^{m_{z}}$, and let $\mathcal{Y}=\langle c_{y},G_{y},A_{y},b_{y}\rangle\subset\mathbb{R}^{n_{y}}$ with $c_{y}\in\mathbb{R}^{n_{y}}$, $G_{y}\in\mathbb{R}^{n_{y}\times\xi_{y}}$, $A_{y}\in\mathbb{R}^{m_{y}\times\xi_{y}}$, and $b_{y}\in\mathbb{R}^{m_{y}}$, then [33] provides the following expression for computing the generalized intersection:
$\mathcal{Z}\cap_{M}\mathcal{Y}=\left\langle c_{z},\left[\begin{array}[]{cc}G_{z}&0_{n_{z}\times\xi_{y}}\end{array}\right],\left[\begin{array}[]{cc}A_{z}&0_{m_{z}\times\xi_{y}}\\ 0_{m_{y}\times\xi_{z}}&A_{y}\\ MG_{z}&-G_{y}\end{array}\right],\left[\begin{array}[]{c}b_{z}\\ b_{y}\\ c_{y}-Mc_{z}\end{array}\right]\right\rangle.$ (2)
2.2.5 Radius of a zonotope
The radius of a zonotope, or a constrained zonotope, $\mathcal{Z}\subset\mathbb{R}^{n}$ is computed as follows
$\mathrm{rad}(\mathcal{Z})=\min\Delta~{}\text{subject to}~{}\mathcal{Z}\subseteq\mathcal{H}(c_{z},\Delta)$
where $\Delta$ is the radius of a minimal hypercube of dimension $n$ denoted by
$\mathcal{H}(c_{z},\Delta)\coloneq\langle c_{z},\Delta I_{n}\rangle$
which is centered at $c_{z}$ and inscribes $\mathcal{Z}$. Notice that for any point $p\in\mathcal{Z}$, it holds that
$\|c_{z}-p\|\leq\sqrt{n}\,\mathrm{rad}(\mathcal{Z})$
where $\mathrm{rad}(\mathcal{Z})=\sup_{p\in\mathcal{Z}}\|c_{z}-p\|_{\infty}$.
3 Problem Definition
Consider a linear time-invariant (LTI) system
$x(k+1)$ $=Ax(k)+Bu(k)+w(k)$ (3a)
$y_{i}(k)$ $=C_{i}x(k)+v_{i}(k)+a_{i}(k),\quad i\in\mathbb{Z}_{[1,p]}$ (3b)
where $x(k)\in\mathbb{R}^{n_{x}}$ is the state, $u(k)\in\mathbb{R}^{n_{u}}$ is a known input, and $y_{i}(k)\in\mathbb{R}^{m_{i}}$ is the measured output of the $i$-th sensor with $i\in\mathbb{Z}_{[1,p]}$. The vector $w(k)\in\mathcal{W}$ represents the process noise, which is bounded and assumed to be contained in the zonotope $\mathcal{W}=\langle c_{w},G_{w}\rangle$, and the vector $v_{i}(k)\in\mathcal{V}_{i}$ represents the measurement noise of the $i$-th sensor, which is also bounded and assumed to be contained in the zonotope $\mathcal{V}_{i}=\langle c_{v_{i}},G_{v_{i}}\rangle$, for every $i\in\mathbb{Z}_{[1,p]}$. Sensor $i$’s measurement at time $k$ may be corrupted by an arbitrary and unbounded attack signal $a_{i}(k)\in\mathbb{R}^{m_{i}}$, which is assumed to be injected by a malicious attacker.
Assumption 1.
We assume the following:
  1. (i)
    Maximum number of attacked sensors: The attacker can attack up to $q\in\mathbb{Z}_{[0,p-1]}$ number of sensors. Although the upper bound $q$ is known, the exact number and/or the set of attacked sensors need not be known.
  2. (ii)
    Redundant observability: There exists $c_{\mathsf{J}}\leq p-q$ such that, for every subset $\mathsf{J}\subset\mathbb{Z}_{[1,p]}$ of sensors with cardinality $|\mathsf{J}|=c_{\mathsf{J}}$, the pair $(A,C_{\mathsf{J}})$ is observable, where $C_{\mathsf{J}}$ is obtained by stacking all $C_{i}$, $i\in\mathsf{J}$, in row blocks.
  3. (iii)
    Knowledge of an initial set: At time $k=0$, we know a compact set $\mathcal{X}_{0}\subset\mathbb{R}^{n_{x}}$ containing the initial state $x(0)\in\mathcal{X}_{0}$.
  4. (iv)
    Bounded noise: The process and measurement noise zonotopes $\mathcal{W},\mathcal{V}_{1},\dots,\mathcal{V}_{p}$ are known.
Assumption 1(i) is fundamental in this paper because it ensures that, at every time $k\in\mathbb{Z}_{\geq 0}$, there exists a set of uncompromised (or safe) sensors:
$\mathsf{S}_{k}\subset\mathbb{Z}_{[1,p]}~{}\text{with}~{}|\mathsf{S}_{k}|\geq p-q$
such that $a_{i}(k)=0_{m_{i}}$ for every $i\in\mathsf{S}_{k}$. This, along with the redundant observability (Assumption 1(ii)), allows us to ensure that the true state can be theoretically reconstructed from the set of uncompromised sensors under the absence of noise. In addition, the assumption entails that the attacker, even though omniscient about the system dynamics and noise bounds, has limited resources at hand and cannot attack all the sensors. We remark that this assumption is not restrictive because it neither restricts the set of attacked sensors to be static with respect to time nor requires that $q$ is less than half the number of sensors $p$ — an assumption that is fundamental in the point-based secure state estimation literature. On the contrary, at any time instant, our problem setup allows the attacker to inject arbitrary signals to any subset of sensors with cardinality less than or equal to $q$, where $q$ is only required to be strictly less than $p$ — meaning that at least one sensor needs to be safe at any time instant.
Assumption 1(ii) is required to enable decentralized set-based operations for secure state estimation without violating the robustness guarantees. Moreover, because Assumption 1(i) allows the attacker to attack up to $q\leq p-1$ sensors, it is necessary that the observability is guaranteed from the remaining safe sensors.
Assumption 1(iii) is required to initialize the set-based state estimation algorithm with $\hat{\mathcal{X}}_{0}=\mathcal{X}_{0}$ such that the inclusion of the state $x(k)$ in the estimated set $\hat{\mathcal{X}}_{k}\subset\mathbb{R}^{n_{x}}$ can be guaranteed. Note that Assumption (iii) does not require any bound on the radius of $\mathcal{X}_{0}$. The only requirement is that the initial set $\mathcal{X}_{0}$ contains $x(0)$. Finally, Assumption 1(iv) is a standard assumption in robust estimation and control literature; see [42] and [15].
Given the uncertain system (3) subject to Assumption 1, we study the following problems:
  1. P1.
    Estimate a set ${\hat{\mathcal{X}}_{k}\subset\mathcal{X}}$ guaranteeing the inclusion $x(k)\in\hat{\mathcal{X}}_{k}$ for every $k\in\mathbb{Z}_{\geq 1}$, where $x(k)$ is the true state of (3).
  2. P2.
    Provide conditions on the magnitude of the attack signals $a_{i}(k)$ such that they can be detected, identified, and/or filtered.
  3. P3.
    Under what conditions, $\hat{\mathcal{X}}_{k}$ satisfies the following stability condition
    $\mathrm{rad}(\hat{\mathcal{X}}_{k})\leq\beta(\mathrm{rad}(\mathcal{X}),k)+\gamma_{1}(\mathrm{rad}(\mathcal{W}))+\gamma_{2}(\mathrm{rad}(\mathcal{V}))$ (4)
    irrespective of the attack signals $a_{i}(k)$, where $\beta$ is a class-$\mathcal{KL}$ function, $\gamma_{1},\gamma_{2}$ are class-$\mathcal{K}_{\infty}$ function [^2] A continuous function $\gamma:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is a class $\mathcal{K}$ function, if it is strictly increasing and $\gamma(0)=0$. It is class $\mathcal{K}_{\infty}$ if it is also unbounded. A continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is a class $\mathcal{KL}$ function, if: (i) $\beta(.,s)$ is a class $\mathcal{K}$ function for any $s\geq 0$; (ii) $\beta(r,.)$ is non-increasing and (iii) $\beta(r,s)\to 0$ as $s\to\infty$ for any $r\geq 0$., and $\mathrm{rad}(\mathcal{V})=\max(\mathrm{rad}(\mathcal{V}_{1}),\dots,\mathrm{rad}(\mathcal{V}_{p}))$.
4 Secure Set-based State Estimation Algorithm
This section presents our secure set-based state estimation (S3E) algorithm. The algorithm comprises four steps that are summarized in Algorithm 1. In the following subsections, we describe each step of the algorithm in detail, provide a comparison with point-based estimators, and address the complexity issue.
4.1 Time Update
The first step in the S3E algorithm is the time update
$\hat{\mathcal{X}}_{k|k-1}=A\hat{\mathcal{X}}_{k-1}\oplus Bu(k-1)\oplus\mathcal{W},$ (5)
where $\mathcal{W}$ is the zonotope that bounds the process noise $w(k)$, and $\hat{\mathcal{X}}_{k-1}$ is the estimated set computed in the previous time step. The computation of $\hat{\mathcal{X}}_{k}$ is described in Section 4.4 with the initialization $\hat{\mathcal{X}}_{0}=\mathcal{X}_{0}$, where $\mathcal{X}_{0}$ is the known initial set from Assumption 1(iii).
Given $x(k-1)\in\hat{\mathcal{X}}_{k-1}$, the time update set $\hat{\mathcal{X}}_{k|k-1}\ni x(k)$ is the set of states to which the system can evolve at time $k\in\mathbb{Z}_{\geq 0}$ subject to the model $(A,B)$, the input $u(k-1)$, and the noise zonotope $\mathcal{W}$. Although the attacker cannot directly influence the time update set as seen from (5), it can do so indirectly through the previous estimate $\hat{\mathcal{X}}_{k-1}$. Thus, the resilience against sensor attacks is achieved by carefully devising the estimated set $\hat{\mathcal{X}}_{k}$.
4.2 Measurement Update
The time update (5) uses model-based information to compute a set $\hat{\mathcal{X}}_{k|k-1}$ containing all points that could be reached by the system state in one time step if they are initialized at the previous estimated set $\hat{\mathcal{X}}_{k-1}$. The measurement update, on the other hand, corrects the conservative estimate of the model-based time update by restricting it according to the sensor measurements.
For every $i\in\mathbb{Z}_{[1,p]}$, let
$\mathcal{Y}_{k}^{i}\coloneq y_{i}(k)-\mathcal{V}_{i}=\langle y_{i}(k)-c_{v_{i}},G_{v_{i}}\rangle$ (6)
be sensor $i$’s output measurement set. Then, the measurement update of sensor $i$ is given by the following generalized intersection of the time update set $\hat{\mathcal{X}}_{k|k-1}$ with the output measurement set $\mathcal{Y}_{k}^{i}$ of sensor $i$ as stated below
$\hat{\mathcal{Z}}_{k}^{i}:=\hat{\mathcal{X}}_{k|k-1}\cap_{C_{i}}\mathcal{Y}_{k}^{i}=\{x\in\hat{\mathcal{X}}_{k|k-1}:C_{i}\hat{\mathcal{X}}_{k|k-1}\in\mathcal{Y}_{k}^{i}\}$ (7)
which can be computed using (2).
The measurement update $\hat{\mathcal{Z}}_{k}^{i}$ is a subset of the time update set $\hat{\mathcal{X}}_{k|k-1}$ that is consistent with the measurement of sensor $i$. That is, given the measurement noise zonotope $\mathcal{V}_{i}$, $\hat{\mathcal{Z}}_{k}^{i}$ is the set of states $x(k)\in\hat{\mathcal{X}}_{k|k-1}$ that could have produced the measurement $y_{i}(k)$. Notice that $\hat{\mathcal{Z}}_{k}^{i}$ could be an empty set if the magnitude of the attack signal $a_{i}(k)$ is sufficiently large.
4.3 Set-based Agreement Protocol
Let $n_{\mathsf{J}}:={p\choose c_{\mathsf{J}}}$, where $c_{\mathsf{J}}\leq p-q$ is the redundant observability parameter. Let $\mathsf{J}_{1},\dots,\mathsf{J}_{n_{\mathsf{J}}}\subset\mathbb{Z}_{[1,p]}$ be the disjoint subsets of size $|\mathsf{J}_{h}|=c_{\mathsf{J}}$, for $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$. The agreement protocol between different measurement updates is given by
$\mathcal{I}_{k}^{h}:=\bigcap_{j\in\mathsf{J}_{h}}\hat{\mathcal{Z}}_{k}^{j}.$ (8)
Theorem 1.
Let Assumption 1 hold and assume $x(k)\in\hat{\mathcal{X}}_{k|k-1}$. Then, at every time $k$, there exists $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$ such that the agreement set $\mathcal{I}_{k}^{h}$ is non-empty and contains the true state $x(k)$.
Proof.
We first show that the measurement update $\hat{\mathcal{Z}}_{k}^{i}$ contains the true state $x(k)$ if sensor $i$ is in the uncompromised set of sensors $\mathsf{S}_{k}$ at time $k$. If $i$ is uncompromised at time $k$, then $a_{i}(k)=0$ and $y_{i}(k)=C_{i}x(k)+v_{i}(k)$. From (6), we have $\mathcal{Y}_{k}^{i}=\langle c_{y_{i}(k)},G_{v_{i}}\rangle$ with
$c_{y_{i}(k)}=y_{i}(k)-c_{v_{i}}=C_{i}x(k)+v_{i}(k)-c_{v_{i}}.$
As $\mathcal{Y}_{k}^{i}$ and the noise zonotope $\mathcal{V}_{i}$ have the same generator matrix $G_{v_{i}}$, we have $\mathrm{rad}(\mathcal{Y}_{k}^{i})=\mathrm{rad}(\mathcal{V}_{i})$. Moreover, it holds that
$\mathcal{Y}_{k}^{i}\supseteq\{c_{y_{i}(k)}-\xi:\|\xi\|_{\infty}\leq\mathrm{rad}(\mathcal{Y}_{k}^{i})\}.$
Thus,
$c_{y_{i}(k)}-v_{i}(k)+c_{v_{i}}=C_{i}x(k)\in\mathcal{Y}_{k}^{i}$
because $\|v_{i}(k)-c_{v_{i}}\|_{\infty}\leq\mathrm{rad}(\mathcal{V}_{i})=\mathrm{rad}(\mathcal{Y}_{k}^{i})$. Since $x(k)\in\hat{\mathcal{X}}_{k|k-1}$ and $C_{i}x(k)\in\mathcal{Y}_{k}^{i}$, it implies that $\hat{\mathcal{Z}}_{k}^{i}$ in (7) is non-empty and contains $x(k)$.
We then show that, at every time $k$, there is a subset $\mathsf{J}_{h}$ of sensors with $|\mathsf{J}_{h}|=c_{\mathsf{J}}$ such that $\mathsf{J}_{h}\subseteq\mathsf{S}_{k}$. This is straightforward because $q$ out of $p$ sensors are compromised and $c_{\mathsf{J}}\leq p-q$, so there must be at least one subset $\mathsf{J}_{h}$ with $|\mathsf{J}_{h}|=c_{\mathsf{J}}$ that contains the uncompromised sensors from the guaranteed $p-q$ uncompromised sensors. In other words, at every time $k\in\mathbb{Z}_{\geq 0}$, there exists $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$ such that $\mathsf{J}_{h}\subseteq\mathsf{S}_{k}$. Since $x(k)\in\hat{\mathcal{Z}}_{k}^{i}$ for every $i\in\mathsf{J}_{h}\subseteq\mathsf{S}_{k}$, the intersection in (8) is non-empty and $x(k)\in\mathcal{I}_{k}^{h}$. ∎
We have shown that the measurement update $\hat{\mathcal{Z}}_{k}^{i}$ of sensor $i$ contains the true state $x(k)$ if it is uncompromised by the attacker at time $k$. Thus, the intersection between the measurement updates of any subset of uncompromised sensors must also contain $x(k)$. In the proof of Theorem 1, we show that by removing the number of attacked sensors $q$, we guarantee the existence of at least one subset $\mathsf{J}_{h}$ with cardinality $|\mathsf{J}_{h}|=c_{\mathsf{J}}$ which is attack-free. Therefore, the corresponding agreement set $\mathcal{I}_{k}^{h}$ must be non-empty and contain $x(k)$.
It is equivalently true to say that if the agreement set $\mathcal{I}_{k}^{h}$ is empty, then there must be at least one sensor in $\mathsf{J}_{h}$ that is compromised by the attacker. Notice that this is only a sufficient condition for detecting an attack because a stealthy attacker can design an attack that ensures that $\mathcal{I}_{k}^{h}$ is non-empty at any time $k$. However, this comes at a cost to the attacker, which is a discussion reserved for later sections in the paper.
4.4 Estimated Set
The estimated set is obtained by taking the union of all agreement sets:
$\hat{\mathcal{X}}_{k}:=\bigcup_{h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}}\mathcal{I}_{k}^{h},\quad k\in\mathbb{Z}_{\geq 1}$ (9)
where $\hat{\mathcal{X}}_{0}=\mathcal{X}_{0}$.
Theorem 2.
Let Assumption 1 hold. Then, for every time $k\in\mathbb{Z}_{\geq 1}$, the inclusion $x(k)\in\hat{\mathcal{X}}_{k}$ is guaranteed, where the estimate $\hat{\mathcal{X}}_{k}$ is computed according to (9).
Proof.
We prove this result by induction. By Assumption 1(iii), we have $x(0)\in\mathcal{X}_{0}=\hat{\mathcal{X}}_{0}$. Because $w(0)\in\mathcal{W}$, the inclusion holds for the time update at $k=1$,
$x(1)\in\hat{\mathcal{X}}_{1|0}=A\mathcal{X}\oplus Bu(0)\oplus\mathcal{W}.$
Therefore, Theorem 1 can be applied, implying the existence of $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$ such that $x(1)\in\mathcal{I}_{1}^{h}$, which gives
$x(1)\in\hat{\mathcal{X}}_{1}=\bigcup_{h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}}\mathcal{I}_{1}^{h}.$
This, in turn, implies that the time update at $k=2$ contains the true state,
$x(2)\in\hat{\mathcal{X}}_{2|1}=A\hat{\mathcal{X}}_{1}\oplus Bu(1)\oplus\mathcal{W}.$
Now, let us assume $x(k^{\prime})\in\hat{\mathcal{X}}_{k^{\prime}|k^{\prime}-1}$ for some $k^{\prime}\in\mathbb{Z}_{\geq 2}$. Then, by Theorem 1, there exists $h\in\mathbb{Z}_{[1,n_{c}]}$ such that $x(k^{\prime})\in\mathcal{I}_{k^{\prime}}^{h}$, implying
$x(k^{\prime})\in\hat{\mathcal{X}}_{k^{\prime}}=\bigcup_{h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}}\mathcal{I}_{k^{\prime}}^{h}.$
Therefore, we have the inclusion at the next time update,
$x(k^{\prime}+1)\in\hat{\mathcal{X}}_{k^{\prime}+1|k^{\prime}}=A\hat{\mathcal{X}}_{k^{\prime}}\oplus Bu(k^{\prime})\oplus\mathcal{W}.$
Thus, the proof is completed because we showed that, for every $k\in\mathbb{Z}_{\geq 1}$, $x(k-1)\in\hat{\mathcal{X}}_{k-1|k-2}$ implies $x(k-1)\in\hat{\mathcal{X}}_{k-1}$, which, in turn, implies $x(k)\in\hat{\mathcal{X}}_{k|k-1}$. Hence, $x(k)\in\hat{\mathcal{X}}_{k}$. ∎
Although the above theorem guarantees the inclusion of the true state, it is important to remark that the number of (constrained) zonotopes in the estimated set $\hat{\mathcal{X}}_{k}$ in (9) may increase with respect to time if the attacker is intelligent and stealthy. We address this issue in Section 4.8 by proposing several techniques that facilitate the computational efficiency of the algorithm. However, it is worth mentioning that the proposed estimation algorithm is resilient because the attacker cannot deteriorate the estimation accuracy arbitrarily over time. If a sensor is injected with a large-valued attack signal, it will be automatically discarded by either yielding an empty measurement update or an empty agreement sets.
The overall algorithm is summarized in Algorithm 1.
Algorithm 1 Secure set-based state estimation
Initialize: $\hat{\mathcal{X}}_{0}=\mathcal{X}_{0}$
for $k=1,2,3,\dots$ do
     Time update (5): $\hat{\mathcal{X}}_{k|k-1}$.
     Measurement update (7): $\hat{\mathcal{Z}}_{k}^{i}$ for every $i\in\mathbb{Z}_{[1,p]}$.
     Agreement protocol (8): $\mathcal{I}_{k}^{h}$ for $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$.
     Estimate (9): $\hat{\mathcal{X}}_{k}$.
end for
4.5 Illustrative Example
We illustrate Algorithm 1 on a simple example of $p=3$ sensors, where $q=1$ sensor has been corrupted. For the purposes of this illustration in Figure 1, we chose all the sensors to have the same dimension $\mathbb{R}^{n_{y}}$ and the first sensor $y_{1}$ to be corrupted. For each sensor $i$, we obtain its corresponding output measurement set $\mathcal{Y}_{k}^{i}$ in the output space $\mathbb{R}^{n_{y}}$ as depicted in Figure 1, which do not necessarily intersect as each sensor $i$ can be measuring different components of the state $x(k)\in\mathbb{R}^{n_{x}}$.
Next, the time update set $\hat{\mathcal{X}}_{k|k-1}$ is constructed according to (5) and shown in the second box from the top in Figure 1. Each output measurement set $\mathcal{Y}_{k}^{i}$ is mapped back to the state space $\mathbb{R}^{n_{x}}$ and constrained to be in the time update set $\hat{\mathcal{X}}_{k|k-1}$ as shown in the third box from the top in Figure 1. The intersection of the measurement update sets $\hat{\mathcal{Z}}^{i}_{k}$ is informative as the intersections of the attack-free $\hat{\mathcal{Z}}^{i}_{k}$ is non-empty. However, the measurement update set $\hat{\mathcal{Z}}^{i}_{k}$ corresponding to a sensor $i$ that is corrupted could also intersect with the other attack-free measurement update sets. This scenario is depicted in Figure 1 where the measurement update set $\hat{\mathcal{Z}}^{1}_{k}$ of the corrupted sensor $1$ intersects with the attack-free measurement update set $\hat{\mathcal{Z}}^{2}_{k}$ due to an intelligently designed attack signal $a_{1}$. We discuss classes of attack signals which can be detected by our algorithm in Section 5.
Since the standing assumption of the algorithm is that the identity of the corrupted sensor(s) is unknown at any given time step $k$, but the number of corrupted sensors $q$ is known, we have to check for non-empty intersections of all $p-q$ combinations of the measurement update sets $\hat{\mathcal{Z}}_{k}^{i}$, which totals to $\binom{p}{p-q}$ checks. Its union then forms our estimated set $\hat{\mathcal{X}}_{k}$ as depicted in the bottom box of Figure 1.
Figure 1: Illustration of Algorithm 1 with $p=3$ and $q=1$. The algorithm deduces that either sensor $1$ or $3$ has been compromised since their corresponding measurement update sets $\hat{\mathcal{Z}}^{1}_{k}$ and $\hat{\mathcal{Z}}^{3}_{k}$ do not intersect.
4.6 Advantage of S3E Algorithm over Point-based Methods
Point-based secure state estimators provide a point-based estimate $\hat{x}(k)\in\mathbb{R}^{n_{x}}$. The security guarantee for these estimators against sensor attacks is an estimation error bound that only depends on the noise and is independent of the attack signal; see [12, 29, 36, 19, 27, 11, 21, 16]. However, for providing such a security guarantee, the fundamental assumption is that the number of attacked sensors is strictly less than half of the total number of sensors. The estimation error bound for point-based estimators has the following form
$\|\hat{x}(k)-x(k)\|\leq\beta(\|\hat{x}(0)-x(0)\|,k)+\gamma_{1}(\|w\|_{\ell^{\infty}})+\gamma_{2}(\|v\|_{\ell^{\infty}})+\gamma_{3}(\|a\|_{\ell^{\infty}})\mathbb{1}_{q\geq\frac{p}{2}}$ (10)
where $\beta$ is a $\mathcal{KL}$ function, $\gamma_{1},\gamma_{2},\gamma_{3}$ are $\mathcal{K}_{\infty}$ functions [^3] A continuous function $\gamma:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is a class $\mathcal{K}$ function, if it is strictly increasing and $\gamma(0)=0$. It is a class $\mathcal{K}_{\infty}$ function if it is also unbounded. A continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is a class $\mathcal{KL}$ function, if: (i) $\beta(.,s)$ is a class $\mathcal{K}$ function for any $s\geq 0$; (ii) $\beta(r,.)$ is non-increasing and (iii) $\beta(r,s)\to 0$ as $s\to\infty$ for any $r\geq 0$., and $\mathbb{1}_{q\geq\frac{p}{2}}$ is the indicator function, which is $1$ when $q\geq\frac{p}{2}$ and 00 otherwise. Here, $v=[\begin{array}[]{ccc}v_{1}^{\top}&\dots&v_{p}^{\top}\end{array}]^{\top}$ and $a=[\begin{array}[]{ccc}a_{1}^{\top}&\dots&a_{p}^{\top}\end{array}]^{\top}$ concatenate the measurement noise and attack signals in all sensors in a column vector, respectively.
It is easy to see that if the number of sensors $p$ is not strictly twice the number of corrupted sensors $q$ (i.e., the assumption of $q<p/2$ is violated), the estimation error bound depends on the magnitude of the attack signal. In this case, the attacker can arbitrarily destroy the accuracy of point-based estimators by injecting attack signals of very large magnitude. Our set-based approach, on the other hand, avoids this situation because it restricts the attack space by only allowing signals of small magnitude. Large attack signals are automatically discarded by either the measurement update (7) or the agreement protocol (8).
Another important advantage of our set-based approach is that the estimated set $\hat{\mathcal{X}}_{k}$ contains a collection of constrained zonotopes, which serve as multiple hypotheses for the true state $x(k)$ in the state space. Although beyond the scope of this paper, an interesting research direction is to developing a hypothesis testing technique to figure out the true set from the sets influenced by the attack signals. For point-based estimators, on the other, the estimation error bound (10) provides a single box of where the state might be. Moreover, the bounds like (10) that are based on comparison functions $\beta\in\mathcal{KL}$ and $\gamma\in\mathcal{K}_{\infty}$ are good for theoretical analysis but are usually very conservative in practice. Therefore, even under no attacks, the value of the right-hand side of (10) is very large as compared to the set based bound presented in the previous section.
4.7 Bound on the Set-based Estimation Error
Since Theorem 2 guarantees that the true state $x(k)$ of system (3) lies in at least one of the constrained zonotopes in the estimated collection of sets $\hat{\mathcal{X}}_{k}$ at each $k\in\mathbb{Z}_{\geq 0}$, it must also lie in a zonotope that overbounds $\hat{\mathcal{X}}_{k}$. That is, let $\hat{\mathcal{X}}_{k}$ be overbounded by a zonotope
$\bar{\mathcal{X}}_{k}=\langle\hat{c}_{x}(k),\hat{G}_{x}(k)\rangle$ (11)
which is obtained by solving
$\min\text{rad}(\bar{\mathcal{X}}_{k})~{}\text{subject to}~{}\hat{\mathcal{X}}_{k}\subseteq\bar{\mathcal{X}}_{k}.$ (12)
Then, the estimation error can be bounded by
$\|\hat{c}_{x}(k)-x(k)\|\leq\text{rad}(\bar{\mathcal{X}}_{k}).$
The stability analysis of this error bound is provided in Section 6 by using a point-based secure estimator. Nonetheless, we remark that this error bound is significantly smaller in practice than the error bounds obtained by point-based secure estimators.
4.8 Methods to Improve Computational Complexity
A major drawback of Algorithm 1 is that the attacker can design intelligent and stealthy attacks such that the number of constrained zonotopes in the estimate $\hat{\mathcal{X}}_{k}$ keep growing with time. Although generating such attacks is a difficult problem as we argue in Section 5, the computational complexity of the algorithm may increase exponentially with time $k$ in the worst case scenario. Therefore, it is important to reduce the complexity by adopting several pruning methods that are described below.
The first obvious step is to remove any empty sets or subsets of other sets in the estimated set $\hat{\mathcal{X}}_{k}$. It is also possible to obtain a single overbounding zonotope [4] of $\hat{\mathcal{X}}_{k}$ as in (12), and use it instead of $\hat{\mathcal{X}}_{k}$ in the next time update. However, a better trade-off between accuracy and complexity is to not overbound the whole collection but only the intersecting zonotopes in the collection $\hat{\mathcal{X}}_{k}$. This may not make the cardinality of $\hat{\mathcal{X}}_{k}$ equal to one, but it reduces it significantly by allowing minimal loss of accuracy.
Another method is to employ zonotope reduction methods [40] to reduce the number of generators in the zonotopes, which are often increased when performing the Minkowski sum operations. However, this technique may result in a larger radius of $\hat{\mathcal{X}}_{k}$ and can also yield conservative estimates.
Finally, one could employ a secure point-based state estimator, if it exists, in parallel with the set-based secure estimator. In this case, we may consider only those candidates in the measurement update collection that lie within the intersection of $\hat{\mathcal{X}}_{k}$ and an error margin generated by a secure point-based state estimator. However, as argued before, the existing point-based estimators require that the total number of sensors be strictly greater than twice the number of compromised sensors $2q<p$ and also some works assume that the members of the attacked sensors also remain unchanged over time, which are tighter requirements than our standing Assumption 1(i). Moreover, the error margins obtained by point-based estimators are usually very conservative. Nonetheless, point-based techniques are useful in the stability analysis of the proposed secure set-based state estimator, as discussed in Section 6.
5 Attack Detection, Identification, and Filtering
In this section, we provide sufficient conditions that guarantee security of the S3E algorithm. First, we provide a condition under which we can detect whether there is an attack on at least one sensor. Second, we provide lower bounds on the magnitude of the attack signals, which, if satisfied, allow us to identify the attacked sensor and/or filter the attack signal to render it ineffective. An attack can be detected whenever the estimated set contains more than one non-intersecting constrained zonotopes. However, an attack is only identified and filtered out either in the measurement update or in the agreement protocol. We also provide a simple attack identification algorithm based on the derived sufficient conditions. The attacks that do not satisfy these conditions are considered to be stealthy, which may be undetectable under the current framework.
5.1 Attack Detection
We say that the attack is detected if the system operator (defender) knows for certain that at least one sensor is compromised at time $k$, i.e., $|\mathsf{S}_{k}|<p$. While attack detection does not necessarily reveal the specific compromised sensor(s), it allows the system operator to raise the alarm and take necessary steps to secure the system. We provide a sufficient condition for attack detection below.
Proposition 1.
Consider the agreement sets $\mathcal{I}_{k}^{h}$, for $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$, given by (8). Then, if
$\bigcap_{h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}}\mathcal{I}_{k}^{h}=\emptyset$ (13)
there is at least one sensor that is compromised at time $k$.
Proof.
We prove the proposition by showing its logical equivalence: ‘no sensor is compromised’ implies ‘(13) does not hold.’ Since for every $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$, the system is redundantly observable through every subset $\mathsf{J}_{h}$ of sensors with $|\mathsf{J}|=c_{\mathsf{J}}$. If no sensor is compromised, then, by Theorem 1, we have that, for every $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$, the agreement set $\mathcal{I}_{k}^{h}$ contains the true state $x(k)$. Therefore, $\mathcal{I}_{k}^{1}\cap\dots\cap\mathcal{I}_{k}^{n_{\mathsf{J}}}\neq\emptyset$. ∎
The above proposition shows that whenever the estimated set contains more than one constrained zonotopes that are disjoint, there has to be at least one sensor that is compromised. Once the attack is detected, the system operator can take necessary measures to secure the system and bring it to safety. If the attacker wants to stay stealthy, they must design attack signals that ensure that all the agreement sets yield a non-empty intersection. In the next subsections, we show that the magnitude of such a stealthy attack has to be sufficiently small.
5.2 Attack Identification
We say that the attack $a_{i}(k)$ is identified if the system operator knows for certain that sensor $i$ is compromised at time $k$, i.e., $i\notin\mathsf{S}_{k}$. The attack identification allows us to automatically discard the measurements of the compromised sensors from the S3E algorithm. We achieve attack identification through a measurement update (7), which involves a generalized intersection between the time update $\hat{\mathcal{X}}_{k|k-1}$ and the output measurement set $\mathcal{Y}_{k}^{i}$.
If sensor $i$ is attacked at time $k$, we derive a sufficient condition on the magnitude of the attack signal $a_{i}(k)$ that would yield an empty measurement update set $\hat{\mathcal{Z}}_{k}^{i}$. This will automatically identify the attack on sensor $i$ and discard its effect from the estimation algorithm. In the next theorem, we use the notation $\bar{\mathcal{X}}_{k|k-1}$ for representing the overbounding zonotope of the time update $\hat{\mathcal{X}}_{k|k-1}$, i.e., $\hat{\mathcal{X}}_{k|k-1}\subseteq\bar{\mathcal{X}}_{k|k-1}$, where
$\bar{\mathcal{X}}_{k|k-1}\coloneq\langle c_{x}(k|k-1),G_{x}(k|k-1)\rangle.$
Theorem 3.
Suppose sensor $i$ is attacked at time $k$. Then, if
$\|a_{i}(k)\|>\sqrt{m_{i}}[\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})+\mathrm{rad}(\mathcal{V}_{i})]-\|C_{i}[x(k)-c_{x}(k|k-1)]+v_{i}(k)-c_{v_{i}}\|$ (14)
then the measurement update $\hat{\mathcal{Z}}_{k}^{i}=\emptyset$. Recall that $m_{i}$ is the dimension of the output $y_{i}$ of system (3).
Proof.
Given a zonotope $C_{i}\bar{\mathcal{X}}_{k|k-1}$ with center $C_{i}c_{x}(k|k-1)$ and radius $\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})$, we have that any point $p\in\mathbb{R}^{m_{i}}$ is outside the $C_{i}\bar{\mathcal{X}}_{k|k-1}$ if $\|C_{i}c_{x}(k|k-1)-p\|>\sqrt{m_{i}}\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})$. Similarly, for $\mathcal{Y}_{k}^{i}$ whose center is $C_{i}x(k)+v_{i}(k)-c_{v_{i}}+a_{i}(k)$ and radius is $\mathrm{rad}(\mathcal{Y}_{k}^{i})$, any $p\in\mathbb{R}^{m_{i}}$ is outside $\mathcal{Y}_{k}^{i}$ if $\|C_{i}x(k)+v_{i}(k)-c_{v_{i}}+a_{i}(k)-p\|>\sqrt{m_{i}}\mathrm{rad}(\mathcal{Y}_{k}^{i})$. Here, the square root term $\sqrt{m_{i}}$ appears because the radius of a zonotope is a maximum norm, which relates to the Euclidean norm as follows: for any vector $p\in\mathbb{R}^{m_{i}}$, $\|p\|\leq\sqrt{m_{i}}\|p\|_{\infty}$.
If the distance between the centers of zonotopes $C_{i}\bar{\mathcal{X}}_{k|k-1}$ and $\mathcal{Y}_{k}^{i}$ is more than $\sqrt{m_{i}}[\mathrm{rad}(\bar{\mathcal{X}}_{k|k-1})+\mathrm{rad}(\mathcal{Y}_{k}^{i})]$, then any point $x\in\bar{\mathcal{X}}_{k|k-1}$ will be such that $C_{i}x\notin\mathcal{Y}_{k}^{i}$. Therefore, in this case, the measurement update (7) is empty. In other words, we have $\hat{\mathcal{Z}}_{k}^{i}=\emptyset$ if
$\|C_{i}c_{x}(k|k-1)-[C_{i}x(k)+v_{i}(k)-c_{v_{i}}+a_{i}(k)]\|>\sqrt{m_{i}}[\mathrm{rad}(\bar{\mathcal{X}}_{k|k-1})+\mathrm{rad}(\mathcal{Y}_{k}^{i})].$
Moreover, we obtain the following due to the triangle inequality
$\|C_{i}[x(k)-c_{x}(k|k-1)]+v_{i}(k)-c_{v_{i}}\|+\|a_{i}(k)\|\geq\|C_{i}c_{x}(k|k-1)-[C_{i}x(k)+v_{i}(k)-c_{v_{i}}+a_{i}(k)]\|.$
Thus, (14) is established by combining the above two inequalities and, from (6), observing that $\mathrm{rad}(\mathcal{Y}_{k}^{i})=\mathrm{rad}(\mathcal{V}_{i})$. ∎
The sufficient condition in Theorem 3 provides a lower bound for the attack signals that can be identified through a measurement update. If the measurement update set $\hat{\mathcal{Z}}_{k}^{i}$ is empty, then it reveals that sensor $i$ is compromised at time $k$. The implication does not hold in the other direction.
If the attacker wants to stay stealthy and effective, it is necessary to violate the lower bound (14) by choosing attack signals of smaller magnitude. If (14) holds, then the attack $a_{i}(k)$ will be automatically identified and filtered out at the measurement update by yielding $\hat{\mathcal{Z}}_{k}^{i}=\emptyset$.
Moreover, notice that the bound (14) depends on the realization of measurement noise $v_{i}(k)$, which is unknown to both the attacker and the defender. One can lower bound (14) by using $\mathrm{rad}(\mathcal{V}_{i})$ and obtain a deterministic bound, which the attacker can violate to generate stealthy attack signals, as follows
$\begin{array}[]{rcl}\|a_{i}(k)\|&>&\sqrt{m_{i}}[\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})+\mathrm{rad}(\mathcal{V}_{i})]-\|C_{i}[x(k)-c_{x}(k|k-1)]+v_{i}(k)-c_{v_{i}}\|\\[5.0pt] &\geq&\sqrt{m_{i}}[\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})+\mathrm{rad}(\mathcal{V}_{i})]-\|C_{i}[x(k)-c_{x}(k|k-1)]\|-\|v_{i}(k)-c_{v_{i}}\|\\[5.0pt] &\geq&\sqrt{m_{i}}[\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})+\mathrm{rad}(\mathcal{V}_{i})]-\|C_{i}[x(k)-c_{x}(k|k-1)]\|-\sqrt{m_{i}}\mathrm{rad}(\mathcal{V}_{i})\\[5.0pt] &=&\sqrt{m_{i}}\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})-\|C_{i}[x(k)-c_{x}(k|k-1)]\|.\end{array}$
Thus, if the attacker would like to stay stealthy by avoiding an empty measurement update set, they must generate an attack $a_{i}(k)$ that satisfies
$\|a_{i}(k)\|\leq\sqrt{m_{i}}\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})-\|C_{i}[x(k)-c_{x}(k|k-1)]\|.$
Finally, notice that satisfying (14) is a sufficient condition for an empty $\hat{\mathcal{Z}}_{k}^{i}$. Therefore, from the attacker’s perspective, violating (14) is a necessary condition for non-empty $\hat{\mathcal{Z}}_{k}^{i}$. This means that if the attack signal violates (14), the measurement update $\hat{\mathcal{Z}}_{k}^{i}$ could still be empty if $C_{i}\hat{\mathcal{X}}_{k|k-1}$ and $\mathcal{Y}_{k}^{i}$ are inconsistent with each other at time $k$.
5.3 Attack Filtering
We say that the attack $a_{i}(k)$ is filtered out if it is detected and discarded from the agreement protocol. To elucidate, let $i\in\mathsf{J}_{h}$, for some $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$. Then, if the attack signal $a_{i}(k)$ is filtered out, then the corresponding agreement set $\mathcal{I}_{k}^{h}$ is empty and does not affect the estimated set $\hat{\mathcal{X}}_{k}$. Although the attack is filtered out whenever it is identified, filtering does not imply identification in general.
In addition to the measurement update, the attacks could also be filtered out via the intersections in the agreement protocol (8), which could yield an empty agreement set $\mathcal{I}_{k}^{h}$ in case of inconsistencies between the measurement updates of the sensors in $\mathsf{J}_{h}$. However, unlike the measurement update, the attacks may not be always identified whenever some of the agreement sets are empty. This is because a compromised sensor $i$ could be a part of multiple agreement sets $\mathcal{I}_{k}^{h}$, for $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$, some of which may turn out to be non-empty. One can identify the attack $a_{i}(k)$ if, for every $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$ such that $i\in\mathsf{J}_{h}$, we have $\mathcal{I}_{k}^{h}=\emptyset$.
To present the next proposition, we introduce some notation. Recall that $\mathsf{S}_{k}\subseteq\mathbb{Z}_{[1,p]}$ is the subset of uncompromised sensors at time $k$. Similarly, let $\mathsf{A}_{k}=\mathbb{Z}_{[1,p]}\setminus\mathsf{S}_{k}$ be the subset of compromised sensors at time $k$. Define
$\mathcal{I}_{k}^{h,\mathsf{A}_{k}}=\bigcap_{j\in\mathsf{J}_{h}\setminus\mathsf{S}_{k}}\hat{\mathcal{Z}}_{k}^{j}$
to be the agreement set of attacked sensors in $\mathsf{J}_{h}$ and let its overbounding zonotope be
¯$\bar{\mathcal{I}}_{k}^{h,\mathsf{A}_{k}}=\langle c_{\mathcal{I}}^{h,\mathsf{A}_{k}}(k),G_{\mathcal{I}}^{h,\mathsf{A}_{k}}(k)\rangle.$
Similarly, define
$\mathcal{I}_{k}^{h,\mathsf{S}_{k}}=\bigcap_{j\in\mathsf{J}_{h}\setminus\mathsf{A}_{k}}\hat{\mathcal{Z}}_{k}^{j}$
be the agreement set of uncompromised sensors in $\mathsf{J}_{h}$ and let its overbounding zonotope be
¯$\bar{\mathcal{I}}_{k}^{h,\mathsf{S}_{k}}=\langle c_{\mathcal{I}}^{h,\mathsf{S}_{k}}(k),G_{\mathcal{I}}^{h,\mathsf{S}_{k}}(k)\rangle.$
Proposition 2.
Let $\mathsf{J}_{h}\subset\mathbb{Z}_{[1,p]}$ be a subset of sensors with $|\mathsf{J}_{h}|=c_{\mathsf{J}}$. Then, if
$\|c_{\mathcal{I}}^{h,\mathsf{A}_{k}}(k)\|>\sqrt{n_{x}}[\mathrm{rad}(\bar{\mathcal{I}}_{k}^{h,\mathsf{S}_{k}})+\mathrm{rad}(\bar{\mathcal{I}}_{k}^{h,\mathsf{A}_{k}})]-\|c_{\mathcal{I}}^{h,\mathsf{S}_{k}}(k)\|$ (15)
then the agreement set $\mathcal{I}_{k}^{h}=\mathcal{I}_{k}^{h,\mathsf{S}_{k}}\cap\mathcal{I}_{k}^{h,\mathsf{A}_{k}}=\emptyset$. Recall that $n_{x}$ is the dimension of the state $x(k)$ of system (3).
Proof.
The proof follows from a similar lines of argument as the proof of Theorem 3. ∎
The interpretation of the inequality in Proposition 2 is not as straightforward as the one in Theorem 3. If the attacker wants to be effective and remain unfiltered, it is necessary for them to devise an attack strategy that violates (15) at every time instant. Otherwise, the attack is discarded due to an empty agreement set. To violate (15), the attacker must take into account the distance between the intersections of the measurement updates of the compromised and uncompromised sensors.
Notice that $c_{\mathcal{I}}^{h,\mathsf{A}_{k}}(k)$, which is the center of the overbounding zonotope ¯$\bar{\mathcal{I}}_{k}^{h,\mathsf{A}_{k}}$, is actually the Chebyshev center of the agreement set $\mathcal{I}_{k}^{h,\mathsf{A}_{k}}$. This agreement set is formed by the intersections of the measurement update sets of the attacked sensors $\mathsf{A}_{k}$. The centers of these measurement update sets are influenced by the corresponding attack signals $a_{j}(k)$. Therefore, to achieve $\mathcal{I}_{k}^{h}\neq\emptyset$, it is necessary that the attacker coordinates the attacks on sensors in such a way that the condition of Proposition 2 is violated.
We remark again that from the attacker’s perspective, the condition of Proposition 2 is necessary for $\mathcal{I}_{k}^{h}\neq\emptyset$, and is not sufficient. This points to several interesting open problems for designing effective attacks. Beyond Theorem 3 and Propositions 1 and 2, our future work will investigate whether a sufficient condition could be derived for the attacker to ensure that both the measurement updates and the agreement sets are non-empty. Moreover, the complexity of generating such attacks is also one of the important questions deferred for our future work.
5.4 Algorithm for Attack Filtering and Identification
A notable contribution of the set-based state estimation scheme in this paper over other secure schemes is Assumption 1(i), which allows the attacker to compromise not only up to $p-1$ sensors at each time instant but also different subsets of sensors at different times. To this end, we remark that we can detect only those compromised sensors that are injected with non-stealthy attack signals, where by non-stealthiness we mean the attack signals that satisfy the conditions of Theorem 3 and Propositions 1 and 2. The attack identification algorithm is fairly simple and can be summarized as in Algorithm 2.
Algorithm 2 Attack filtering and identification
1:Measurement updates $\hat{\mathcal{Z}}_{k}^{i}$ for $i\in\mathbb{Z}_{[1,p]}$, sensor subsets $\mathsf{J}_{h}$, for $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$, and their corresponding agreement sets $\mathcal{I}_{k}^{h}$.
2:Initialize $\hat{\mathsf{S}}_{k}=\emptyset$
3:$\hat{\mathsf{A}}_{k}\leftarrow\{i\in\mathbb{Z}_{[1,p]}:\hat{\mathcal{Z}}_{k}^{i}=\emptyset\}$
4:for $h=1,\dots,n_{\mathsf{J}}$ do
5:     if $\mathcal{I}_{k}^{h}\neq\emptyset$ then
6:         Estimated safe subset $\hat{\mathsf{S}}_{k}\leftarrow\hat{\mathsf{S}}_{k}\cup\mathsf{J}_{h}$
7:     end if
8:end for
9:Identified attacked sensors $\hat{\mathsf{A}}_{k}\leftarrow\hat{\mathsf{A}}_{k}\cup\{\mathbb{Z}_{[1,p]}\setminus\hat{\mathsf{S}}_{k}\}$.
At time $k$, Algorithm 2 first identifies some of the attacked sensors by checking whether the corresponding measurement update is empty. Then, for each $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$, if the agreement set $\mathcal{I}_{k}^{h}$ is non-empty, then the sensors with indices in $\mathsf{J}_{h}$ are either safe or compromised with a stealthy attack signal. Otherwise, the set $\mathsf{J}_{h}$ contains at least one attacked sensor. By checking all the combinations $\mathsf{J}_{h}$ and storing a ‘potentially’ safe subset of sensors in $\hat{\mathsf{S}}_{k}$ at every iteration, a subset of attacked sensors $\hat{\mathsf{A}}_{k}$ are updated by adding those sensors that are not in $\hat{\mathsf{S}}_{k}$.
Remark 1.
The estimated safe subset $\hat{\mathsf{S}}_{k}$ contains the true safe subset $\mathsf{S}_{k}$ at time $k$, i.e., $\mathsf{S}_{k}\subseteq\hat{\mathsf{S}}_{k}$. If there are sensors that are injected with small-valued stealthy attack signals, our detection algorithm fails to recognize those attacks and considers those sensors to be uncompromised. Therefore, the estimated set of identified attacked sensors $\hat{\mathsf{A}}_{k}$ is only a subset of the true set of attacked sensors $\mathsf{A}_{k}=\mathbb{Z}_{[1,p]}\setminus\mathsf{S}_{k}$.
Remark 2.
In Algorithm 2, we identify a subset of attacked sensors at every time $k$. However, in the time-invariant attack setting where the attacker does not change the subset of sensors to compromise at every time instant $k$, Algorithm 2 can be adapted to cumulatively detect and remove the attacked sensors over time. This can have application in sensor fault detection and isolation as faults can be considered as naive, time-invariant attacks.
6 Stability of the Estimated Set
To assist in the stability analysis of the secure zonotopic state estimation algorithm, in the sense that the estimated set $\hat{\mathcal{X}}_{k}$ containing the true state $x(k)$ is bounded with respect to time $k\in\mathbb{Z}_{\geq 0}$ given that the initial set $\mathcal{X}$ is bounded, we employ a secure point-based state estimator in parallel. Such an estimator provides a point-based estimate of the state and ensures that the bound on the estimation error is unaffected by the attack signal. However, for this section, we restrict Assumption 1(i) by considering $2q<p$. In the following, we define a point-based secure estimator and use it to slightly modify our proposed S3E algorithm, which ensures that the set-based estimate is stable.
6.1 Existence of a Secure Point-based State Estimator
To guarantee stability of S3E algorithm, we employ a point-based estimator which is assumed to be secure with respect to the attack signals $a_{i}$ and robust with respect to disturbances $w$ and noise $v$ as follows.
Assumption 2.
There exists a secure point-based state estimator $E:\mathbb{R}^{m}\times\mathcal{U}\rightarrow\mathbb{R}^{n_{x}}$ which provides a point-based state estimate $\hat{x}(k)=E(y_{1}(k),\dots,y_{p}(k),u(k))$ to system (3) satisfying the following for $k\in\mathbb{Z}_{\geq 0}$,
$\|\hat{x}(k)-x(k)\|\leq\beta(\|\hat{x}(0)-x(0)\|,k)+{\alpha\max(\bar{w},\bar{v})}$ (16)
for some class $\mathcal{KL}$ function $\beta$ and $\alpha\in\mathbb{R}_{\geq 0}$, where $\|w(k)\|_{\infty}\leq\bar{w}$ and $\|v(k)\|_{\infty}\leq\bar{v}$ for all $k\in\mathbb{Z}_{\geq 0}$.
Note that in Assumption 2, the upper bound on the state estimation error is not affected by the attack signals $a_{i}$. We call estimators that possess this property secure and results exist in the literature for linear systems (e.g., [12, 27, 21, 16, 9]). So far, these results require at most $q\leq\lceil p/2\rceil-1$ number of attacked sensors, i.e., strictly less than half the number of sensors can be attacked. Thereby, this imposes a stricter requirement than Assumption 1(i), which we state below.
Assumption 3.
The number of attacked sensors $q\in\mathbb{Z}_{\geq 0}$ satisfies $2q<p$, where $p\in\mathbb{Z}_{\geq 0}$ is the total number of sensors. The integers $p$ and $q$ are known, but the exact attacked sensors are unknown.
In the following subsection, we provide a stability analysis of S3E algorithm by using a secure point-based state estimator in parallel.
6.2 Modified Estimate
We modify the estimate as
$\hat{\mathcal{X}}^{\mathrm{mod}}_{k}=\hat{\mathcal{X}}_{k}\cap\mathcal{H}(\hat{x}(k),\Delta_{k})$ (17)
where $\hat{\mathcal{X}}_{k}$ is the estimate in (9) and $\mathcal{H}(\hat{x}(k),\Delta_{k})$ denotes a hypercube with center $\hat{x}(k)$, which is the estimate provided by a secure point-based estimator under Assumption 2, and radius
$\Delta_{k}\coloneq\beta(\|x(0)-\hat{x}(0)\|,k)+\alpha\max(\bar{w},\bar{v})$ (18)
where $\beta\in\mathcal{KL}$, $\alpha\in\mathbb{R}_{\geq 0}$, $\bar{w}\in\mathbb{R}_{\geq 0}$ and $\bar{v}\in\mathbb{R}_{\geq 0}$ come from Assumption 2. Figure 2 illustrates the modified measurement update (17).
Figure 2: Modified measurement update
It is important to note that the practicality of this pruning algorithm relies on reasonable estimates of $\beta\in\mathcal{KL}$ and $\alpha>0$, which is usually difficult to obtain. For linear systems, they can be obtained easily; however, they are usually quite conservative because of their dependence on the condition number of matrices involved in the estimation. Nonetheless, this technique is useful in providing a theoretical bound of the estimation error.
To this end, we define the estimation error in terms of the center of
$\bar{\mathcal{X}}^{\mathrm{mod}}_{k}:=\langle\hat{c}^{\mathrm{mod}}_{x}(k),\hat{G}^{\mathrm{mod}}_{x}(k)\rangle$ (19)
a zonotope with minimal radius such that it overbounds $\hat{\mathcal{X}}^{\mathrm{mod}}_{k}$ in (17), i.e.,
$\begin{array}[]{rl}\min&\mathrm{rad}(\bar{\mathcal{X}}^{\mathrm{mod}}_{k})\\ \textrm{ subject to }&\hat{\mathcal{X}}^{\mathrm{mod}}_{k}\subseteq\bar{\mathcal{X}}^{\mathrm{mod}}_{k}.\end{array}$ (20)
Since $\lim_{k\rightarrow\infty}\Delta_{k}=\alpha\max(\bar{w},\bar{v})$,
$\lim_{k\rightarrow\infty}\mathrm{rad}(\bar{\mathcal{X}}^{\mathrm{mod}}_{k})\leq\alpha\max(\bar{w},\bar{v}),$
which only depends on the bounds on the process and measurement noise, and not on the attack signals.
6.3 Revised Bound on the Estimation Error
We provide a bound for the estimation error $\hat{c}^{\mathrm{mod}}_{x}(k)-x(k)$, where $\hat{c}^{\mathrm{mod}}_{x}(k)$ is the center of $\bar{\mathcal{X}}^{\mathrm{mod}}_{k}$ in (19).
Proposition 3.
Let Assumptions 1(ii)-(iii), 2, and 3 hold. Consider Algorithm 1 with a modified estimate (17). Then, for every $k\in\mathbb{Z}_{\geq 0}$, the overall estimation error bound is given by the following
$\|\hat{c}^{\mathrm{mod}}_{x}(k)-x(k)\|\leq\Delta_{k}$ (21)
where $\Delta_{k}$ is defined in (18) and $\hat{c}^{\mathrm{mod}}_{x}(k)$ is the center of $\bar{\mathcal{X}}^{\mathrm{mod}}_{k}$ computed by (20).
Proof.
Note that $x(k)\in\hat{\mathcal{X}}_{k}$ according to Theorem 2 and $x(k)\in\mathcal{H}(\hat{x}(k),\Delta_{k})$ according to Assumption 2. Therefore, we have that $x(k)\in\hat{\mathcal{X}}^{\mathrm{mod}}_{k}$. Since $\hat{\mathcal{X}}^{\mathrm{mod}}_{k}\subseteq\bar{\mathcal{X}}^{\mathrm{mod}}_{k}$ because of (20), we obtain (21). ∎
Proposition 3 enables us to conclude that by implementing a secure point-based estimator with properties given by Assumption 2, we can guarantee that the estimation error satisfies (21) with a bound $\Delta_{k}$ that asymptotically converges to a bound that only depends on the bounds on the process and measurement noise, and does not depend on the attack signals $a_{i}(k)$. This is a guarantee that is absent in the purely secure set-based estimation algorithm (Algorithm 1) developed in the previous section. However, we now have a more restrictive Assumption 3 (in comparison to Assumption 1(i)) that guarantees the existence of a secure point-based estimator possessing the property in Assumption 2.
7 Numerical Simulation
We evaluated our proposed algorithms on two examples. In the first illustrative example, we assume that half of the total number of sensors can be targeted by an attacker at every time instant. We illustrate the estimation algorithm and use a strategy to reduce the complexity with a minimal loss of estimation accuracy. In the second example of a three-story building structure, we assume redundant observability from every subset of two sensors. We use the CORA toolbox [4] to generate our simulations.
(a) Time $k=1$. Sensors 2 and 3 are attacked.
(b) Time $k=2$. Sensors 3 and 4 are attacked.
(c) Time $k=3$. Sensors 1 and 4 are attacked.
Figure 3: Snapshots of estimated sets using Algorithm 1. The attack is generated according to (22) where different sensors are attacked at different time steps.
(a) Time $k=1$. Sensors 2 and 3 are attacked.
(b) Time $k=2$. Sensors 3 and 4 are attacked.
(c) Time $k=3$. Sensors 1 and 4 are attacked.
Figure 4: Snapshots of estimated sets using Algorithm 1 while applying a complexity reduction technique to overapproximate the estimated intersecting sets by one set. The attack is generated according to (22) where different sensors are attacked at different time steps.
7.1 Two Dimensional Linear System
In this example, we consider a simple two-dimensional linear system for illustrative purposes. We consider a zero control input $u(k)\equiv 0$, and the system matrices are given by
$A=$ $\left[\begin{array}[]{cc}1&0\\ 1&1\end{array}\right],\quad C_{1}=\left[\begin{array}[]{cc}1&0\\ 0&1\end{array}\right],\quad C_{2}=\left[\begin{array}[]{cc}1&1\\ 1&0\end{array}\right],$
$C_{3}=$ $\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right],\quad C_{4}=\left[\begin{array}[]{cc}1&2\\ 2&1\end{array}\right].$
The number of sensors is $p=4$, and we suppose that an attacker can target any combination of $q=2$ sensors at different time instants. In this case, since $q=p/2$, the point-based state estimators requiring $q<p/2$ cannot be employed.
Let the process noise bound be $\mathcal{W}=\langle 0,\sigma_{\mathcal{W}}I_{2}\rangle$ with $\sigma_{\mathcal{W}}=0.02$ and the measurement noise bounds of the four sensors be $\mathcal{V}_{1}=\mathcal{V}_{2}=\mathcal{V}_{3}=\mathcal{V}_{4}=\langle 0,\sigma_{\mathcal{V}}I_{2}\rangle$ with $\sigma_{\mathcal{V}}=1$, where $I_{2}$ is the identity matrix of dimension $2\times 2$. The attack is generated according to the following relation
$a_{i}(k)\sim U(-\sigma_{\mathcal{V}}\phi(k),\sigma_{\mathcal{V}}\phi(k))$ (22)
where $U$ is a uniform distribution over an interval $(-\sigma_{\mathcal{V}}\phi(k),\sigma_{\mathcal{V}}\phi(k))$ and $\phi(k)$ is strictly increasing sequence with $\phi(1)=1$. The index of the two attacked sensors rotates among the available sensor indices.
The time update, measurement update, agreement, and estimated sets can be computed using Algorithm 1. Fig. 3 illustrates the time update set $\hat{\mathcal{X}}_{k|k-1}$ (green), the measurement sets $\hat{\mathcal{Z}}_{k}^{i}$ (pink (unattacked) and red (attacked)), the estimated set $\hat{\mathcal{X}}_{k}$ (black), and the true state $x(k)$ (blue) for different sets of sensors attacked at different times. The attacker chooses a random combination of two sensors to compromise at each time instant. At time $k=1$, we have in Fig. 3(a) sensors $2$ and $3$ under attack. Then, sensors $3$ and $4$ are attacked in Fig. 3(b) at time step $k=2$. Notice that the true state stays included in the estimated set $\hat{\mathcal{X}}_{k}$ for every $k$.
The generated attack values lead to an increase in the number of generated sets, as depicted in Fig. 3(c). To manage this complexity, we employ a reduction technique that involves taking the union of intersecting estimated sets. The impact of this technique is illustrated in Fig. 4, where we use the same seed for generating random attacks from (22). This approach can be seen by comparing Fig. 3(a) with Fig. 4(a). The reduced estimated sets are then carried forward to subsequent steps, as shown in Fig. 4(b) and Fig. 4(c). This is why the sets in these figures show slight differences when compared to those in Fig. 3(b) and Fig. 3(c).
(a) Time $k=1$, Sensor 2 attacked.
(b) Time $k=2$, Sensor 3 attacked.
(c) Time $k=3$, Sensor 1 attacked.
Figure 5: Snapshots of estimated sets using Algorithm 1 under time-varying attack, where different sensors are attacked at different time steps.
7.2 Three-story Building Structure
We now consider a three-story building structure of [38] described by a mechanical system
$M\ddot{q}(t)+D\dot{q}(t)+Sq(t)=Gu(t),$ (23)
where $q(t)\in\mathbb{R}^{3}$ is the vector of relative horizontal displacements of the floors and $u(t)\in\mathbb{R}$ is the ground acceleration due to earthquake which is a measured input signal. Also, $M\in\mathbb{R}^{3\times 3}$ is the mass matrix, $D\in\mathbb{R}^{3\times 3}$ is the damping matrix, $S\in\mathbb{R}^{3\times 3}$ is the stiffness matrix, and $G\in\mathbb{R}^{3}$ is the loading vector. The parameter values of the system (23) are provided by [38] as:
[…]
By considering the state $x(t)=[\begin{array}[]{cc}q(t)^{\top}&\dot{q}(t)^{\top}\end{array}]^{\top}$, we can obtain the state-space representation in continuous time
$\dot{x}=A_{c}x+B_{c}u$
where
$A_{c}=\left[\begin{array}[]{cc}0_{3\times 3}&I_{3}\\ -M^{-1}S&-M^{-1}D\end{array}\right],\quad B_{c}=\left[\begin{array}[]{c}0_{3\times 1}\\ -M^{-1}G\end{array}\right].$
After discretization with sample time $\delta=10^{-3}$, we obtain the system in the form (3a), where
$A$ $=\exp(A_{c}\delta)$
$B$ $=A_{c}^{-1}(A-I_{6})B_{c}.$
Here, our goal is to monitor the building dynamics under an earthquake using secure set-based state estimation algorithm.
We assume that each floor of the building is equipped with a sensor, i.e., $p=3$, that measures the relative displacement and the velocity of that floor, which can be collected in the output vector $y_{i}(k)\in\mathbb{R}^{3}$ as given by (3b), for $i\in\mathbb{Z}_{[1,3]}$, where
$\begin{array}[]{c}C_{1}=\left[\begin{array}[]{cccccc}1&-1&0&0&0&0\\ 1&0&-1&0&0&0\\ 0&0&0&1&0&0\end{array}\right],\quad C_{2}=\left[\begin{array}[]{cccccc}-1&1&0&0&0&0\\ 0&1&-1&0&0&0\\ 0&0&0&0&1&0\end{array}\right]\\[21.33955pt] C_{3}=\left[\begin{array}[]{cccccc}-1&0&1&0&0&0\\ 0&-1&1&0&0&0\\ 0&0&0&0&0&1\end{array}\right].\end{array}$
We suppose that the attacker can compromise only one sensor at each time, i.e., $q=1$, whose rationale is explained below. Although the pairs $(A,C_{1})$ ,$(A,C_{2})$, and $(A,C_{3})$ are ‘theoretically’ observable, they are not ‘practically’ observable because the corresponding observability matrices have some singular values very close to zero. Therefore, in this example, we assume redundant observability from every pair of two sensors. Our goal is to monitor the building’s floor displacements and velocities irrespective of the compromised sensor.
Suppose the dynamics are corrupted by the process noise $w(k)$, which is bounded by $\mathcal{W}=\langle 0,\sigma_{\mathcal{W}}I_{6}\rangle$ with $\sigma_{\mathcal{W}}=0.02$. Similarly, the sensor measurements are corrupted by noise $v_{1}(k)$, $v_{2}(k)$, and $v_{3}(k)$, respectively, which are bounded by $\mathcal{V}_{i}=\langle 0,\sigma_{\mathcal{V}}I_{3}\rangle$ with $\sigma_{\mathcal{V}}=1$.
To illustrate the efficacy of our algorithm, we apply a similar attack as in (22) to the sensors in which the attacker randomly chooses a sensor $i$ at every time step $k$ and injects false data into its measurement $y_{i}(k)$. In Fig. 5(a), sensor $2$ is compromised. We compute the estimated set (black), which contains the true state. Then, sensor $3$ is attacked in Fig. 5(b) in which the time update set $\hat{\mathcal{X}}_{k|k-1}$ shrinks due to the progressive intersection. Finally, sensor $1$ is attacked in Fig. 5(c). The true state $x(k)$ remains enclosed by the estimated measurement update sets at every time step. Also, the estimation error remains bounded, and the attacker cannot destroy the accuracy of the set-based state estimate. Finally, we remark that the intersections depicted in the figure may not appear accurate because the figure shows a projection of six-dimensional sets on a two-dimensional plane.
8 Conclusion and future work
We presented a secure set-based state estimation algorithm that can ensure the inclusion of the true state of an LTI system in the estimated set even when all but one sensor could be compromised by an adversarial attacker. Our proposed algorithm overcomes the limitation of point-based secure estimators that restrict the number of attacked sensors to strictly less than half of the total number of sensors. We achieved this by constructing agreement sets from the intersection of various observable combinations of measurement update sets. We showed that our algorithm guarantees the inclusion of the true state in the estimated set. Moreover, we proposed sufficient conditions for detecting, identifying, and filtering the attack signals and presented a simple algorithm to identify the set of compromised sensors at every time instant. The proposed algorithm may find applications in safety verification of safety-critical systems whose multiple sensors could be compromised by an attacker.
While our algorithm’s worst-case complexity may increase with time under intelligently-designed stealthy attacks, we argued that it is challenging for attackers to execute such attacks due to the requirement of a complete understanding of the system and estimation algorithm and substantial computational resources. Nonetheless, we suggested various strategies to reduce the complexity of our algorithm to facilitate its implementation.
We also incorporated a secure point-based state estimator into our algorithm to prune the candidate sets when less than half of the sensors can be compromised. This strategy’s effectiveness depends on an accurate approximation of the guaranteed estimation error provided by the secure point-based state estimator. Nonetheless, the modified algorithm provides asymptotic convergence guarantees with an explicit upper bound that depends only the known process and measurement noise, and not on the attack signals.
Our future work will focus on the set-based secure state estimation of nonlinear systems and developing a data-driven approach for secure estimation when the system model is unknown. We also highlight several open questions that remain unaddressed in this paper. For instance, devising better agreement protocols that can improve the lower bounds on the detectable, identifiable, and filterable attack signals will further strengthen the set-based estimation approach presented in this paper. The question of computational complexity under stealthy attacks remains to be addressed in a rigorous manner. The stability of the estimated set is only addressed for the case when strictly less than half of the sensors are attacked. When more than half of the sensors are attacked, it is challenging to upper bound the estimated set by a contracting set. Finally, from the attacker’s perspective, analyzing the complexity of generating stealthy attacks that result in the increased complexity of the set-based estimation is an interesting research problem.